# A Data-driven Typology of Vision Models from Integrated Representational Metrics

### v2

@@ Abstract
Large vision models differ widely in architecture and training paradigm, yet we lack principled methods to determine which aspects of their representations are shared across families and which reflect distinctive computational strategies. We leverage a suite of representational similarity metrics, each capturing a different facet—geometry, unit tuning, or linear decodability—and assess family separability using multiple complementary measures. Metrics preserving geometry or tuning (e.g., RSA, Soft Matching) yield strong family discrimination, whereas flexible mappings such as Linear Predictivity show weaker separation. These findings indicate that geometry and tuning carry family-specific signatures, while linearly decodable information is more broadly shared. To integrate these complementary facets, we adapt Similarity Network Fusion (SNF), a method inspired by multi-omics integration. SNF achieves substantially sharper family separation than any individual metric and produces robust composite signatures. Clustering of the fused similarity matrix recovers both expected and surprising patterns: supervised ResNets and ViTs form distinct clusters, yet all self-supervised models group together across architectural boundaries. Hybrid architectures (ConvNeXt, Swin) cluster with masked autoencoders, suggesting convergence between architectural modernization and reconstruction-based training. This biology-inspired framework provides a principled typology of vision models, showing that emergent computational strategies—shaped jointly by architecture and training objective—define representational structure beyond surface design categories.
@@ 1 Introduction
The rapid proliferation of vision models—spanning diverse architectures from CNNs to Vision Transformers, and training paradigms from supervised to self-supervised learning—has created a rich ecosystem of computational approaches to visual processing. Yet we lack principled methods to understand which aspects of learned representations are universally shared across this diverse landscape and which are specific signatures of particular model families. Do all vision models converge on similar geometric organizations of their representation spaces? Is linearly accessible information a common currency across architectures, or do different model families encode information in fundamentally distinct ways? These questions are critical for understanding the computational principles underlying successful vision models and for predicting how different models will behave on novel tasks.
Current approaches to model comparison rely heavily on individual similarity metrics—Representational Similarity Analysis (RSA) (Kriegeskorte et al., ), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), Linear Predictivity (Yamins et al., 2014), and others—each capturing a specific facet of representational structure. However, this fragmented methodology obscures a crucial insight: different representational facets (geometry, tuning properties, linearly accessible information) may vary in their universality across model families. Some aspects might reflect convergent computational solutions shared broadly across architectures, while others might constitute distinctive signatures of specific model families. Understanding this landscape requires not just comparing models, but systematically evaluating which representational properties are conserved versus specialized. Furthermore, relying on a single representational similarity metric is inherently fragile and can easily enable cherry-picking. A growing body of work shows that commonly used similarity metrics can yield contradictory views of model relatedness: tuning-sensitive metrics (e.g., SoftMatch) distinguish architectural families, whereas geometry-preserving measures (e.g., CKA/RSA) are more sensitive to training vs. random initialization but often conflate architectures(Bo et al., 2024). Because each metric emphasizes different invariances, no single metric provides a complete characterization, making conclusions highly dependent on metric choice. This motivates the need for a principled fusion method that integrates complementary representational signals to obtain a more stable and robust organization of model space. In this work, we introduce a framework that addresses these challenges through two key contributions. First, we systematically evaluate how different representational facets discriminate between model families, revealing that geometry or tuning-preserving metrics (RSA, Soft Matching (Khosla & Williams, 2024)) strongly separate families while metrics capturing linearly accessible information show weaker discrimination. This pattern indicates that tuning of individual neurons is family-specific and that linearly accessible signals vary less across these model families. Second, inspired by multi-omics approaches in biology where diverse molecular signatures are integrated to reveal cell types, we employ Similarity Network Fusion (SNF) (Wang et al., 2014) to combine these complementary perspectives into unified model signatures that provide clearer family identity than any single metric alone.
This integrated approach enables us to construct a data-driven typology of vision models—a novel contribution that moves beyond surface-level architectural categories to reveal how models actually organize information. Our typology does not rely on a priori assumptions about which models should group together based on architecture or training method. Instead, following empirical traditions in psychology, neuroscience, and genetics (Wang et al., 2014; Letwin et al., 2006; Echtermeyer et al., 2011; Mukamel & Ngai, 2019) where researchers identify clusters of individuals through correlations across multiple behavioral or molecular indices, we discover natural groupings based on how models process visual information. While typologies could alternatively be defined from theoretical perspectives emphasizing explicit model properties like architecture or training data, our empirical approach reveals surprising patterns: all self-supervised models form a unified cluster that transcends architectural boundaries, with self-supervised ResNets grouping more closely with self-supervised ViTs than with their supervised architectural siblings. Similarly, hybrid architectures (ConvNeXt, Swin) (Liu et al., 2022; 2021) cluster with MAE models, suggesting that architectural modernization and masked reconstruction converge on similar computational strategies despite different design origins.
This data-driven typology provides researchers with a reference framework for understanding where any model instance sits within the broader landscape of vision models. Just as biological taxonomies help scientists understand relationships between species based on genetic and phenotypic traits, our representational typology reveals the “species” of vision models based on their computational strategies. By developing methods to systematically integrate different facets of representation in comparative analyses of models, we provide practical tools for navigating the expanding universe of vision models—enabling researchers to understand model relationships, predict transfer learning compatibility, and make informed choices about which models will exhibit similar behaviors on novel tasks.
@@
Figure 1: Top left: Each representational metric defines a pairwise similarity matrix over models. Bottom left: Each matrix is visualized as an affinity graph, with nodes representing models and edge widths reflecting pairwise similarity strength; weak similarities below a threshold are omitted for clarity. Right: A consensus matrix obtained via Similarity Network Fusion (SNF) highlights relations consistently supported across metrics while leveraging complementary signals. In the fused graph, solid edges denote agreement across all metrics, dotted edges indicate partial support; strong but uncorroborated edges may persist with reduced weight (e.g., edge 4–5); weak AND metric-specific connections are typically suppressed (e.g., edge 1–5).
@@ 2.1 Model Selection and Dataset
We define the model family as the combination of training paradigm and architecture. We analyze 35 vision models across four primary categories: supervised Convolutional Neural Networks (CNNs), self-supervised CNNs, supervised Transformers, and self-supervised Transformers. We treat ConvNeXt (Liu et al., 2022) and Swin (Liu et al., 2021) as distinct families due to their hybrid nature—ConvNeXt incorporates Transformer-inspired design principles within a convolutional architecture, while Swin introduces CNN-like inductive biases into the Transformer framework. We deliberately restricted the model set to encoders pre-trained on ImageNet-1k. This design isolates the effects of architecture and training objective within a shared data and label space, so that differences in training data distribution do not become a dominant confound. For datasets, we use the ImageNet-1k (Deng et al., 2009) and Ecoset (Mehrer et al., 2021) validation sets and the CIFAR10 (Krizhevsky, 2009) and CIFAR100 (Krizhevsky, 2009) test sets. Complete dataset and model details are provided in Appendix A and B.
@@ 2.2 Representational Metrics
We evaluate widely used similarity metrics that differ in the flexibility of the mappings they permit—from permutation-based alignments (soft-matching) to rigid geometric transformations (Procrustes) to looser linear mappings (linear predictivity) as well as non-fitting approaches that compare representational geometry directly (RSA). Consider two representations $\mathbf{X}_{i}\in\mathbb{R}^{M\times N_{i}}$ and $\mathbf{X}_{j}\in\mathbb{R}^{M\times N_{j}}$ from different models, where $M$ denotes the number of stimuli and $N_{i},N_{j}$ denote the number of units. All representations are mean-centered along the sample dimension as required. For metrics requiring a fitting procedure (e.g., Soft matching, linear predictivity, Procrustes), similarity values reflect the mean 5-fold cross-validation score.
Singular Vector Canonical Correlation Analysis (Raghu et al., 2017).
@@
SVCCA first applies singular value decomposition (SVD) to the representation matrices from two models or layers to isolate their most informative directions: $\mathbf{X}_{i}=\mathbf{U}_{i}\Sigma_{i}\mathbf{V}_{i}^{\top}$ and $\mathbf{X}_{j}=\mathbf{U}_{j}\Sigma_{j}\mathbf{V}_{j}^{\top}$. Retaining the top $N^{\prime}_{i}$ and $N^{\prime}_{j}$ singular vectors that explain $99\%$ of the variance yields the reduced representations: $\mathbf{X}^{\prime}_{i}=\mathbf{U}_{i}^{(N^{\prime}_{i})\!\top}\mathbf{X}_{i}$ and $\mathbf{X}^{\prime}_{j}=\mathbf{U}_{j}^{(N^{\prime}_{j})\!\top}\mathbf{X}_{j}$ whose dominant singular directions capture a disproportionate share of the total information. Canonical correlation analysis (CCA) (Hardoon et al., 2004) is then applied to these reduced matrices to find linear projections $\mathbf{A}$ and $\mathbf{B}$ that maximize their correlation: $\mathbf{Q}=\max_{\mathbf{A},\,\mathbf{B}}\operatorname{corr}\!\bigl(\mathbf{A}\mathbf{X}^{\prime}_{i},\;\mathbf{B}\mathbf{X}^{\prime}_{j}\bigr)$ subject to unit-variance constraints, providing a final similarity score that reflects how closely the informative subspaces of the two representations align.
Projection-Weighted Canonical Correlation Analysis (Morcos et al., 2018).
@@
Compared to SVCCA, PWCCA did not apply SVD before CCA and it re-weights the canonical directions according to their contribution to the original representation. After CCA, we obtain canonical vectors $\mathbf{A}$ and $\mathbf{B}$ and their corresponding correlations $q_{i}$ for i=1,2,ki=1,2,...k, where $k=\min(N_{i},N_{j})$. Instead of giving each direction equal weight, PWCCA projects the unreduced representation $\mathbf{X}_{i}$ onto its own canonical vectors to measure how strongly each one reconstructs the data. The projection weight for the $i^{th}$ canonical direction is $\alpha_{i}=\frac{\left\|\mathbf{X}_{i}{\bm{a}}_{i}\right\|_{1}}{\sum_{j=1}^{k}\left\|\mathbf{X}_{i}{\bm{a}}_{j}\right\|_{1}}$ where ${\bm{a}}_{i}$ is the $i^{th}$ column of $\mathbf{A}$. The final PWCCA similarity is then the weighted sum of canonical correlations: $\Sigma_{i=1}^{k}\alpha_{i}q_{i}$ which emphasizes directions that explain the greatest fraction of variance in $X_{i}$ and down-weights noisy, low-variance components, yielding a more faithful measure of representational similarity than the equal-weight SVCCA score.
Linear Centered Kernel Alignment (Kornblith et al., 2019).
@@
Linear CKA provides a scalar measure of how similarly two sets of representations capture the relationships among the same collection of samples. It is is defined as $\frac{\|\mathbf{X}_{i}^{T}\mathbf{X}_{j}\|_{F}^{2}}{\|\mathbf{X}_{i}^{T}\mathbf{X}_{i}\|_{F}\|\mathbf{X}_{j}^{T}\mathbf{X}_{j}\|_{F}}$, where $\|.\|_{F}$ is the frobenius norm, and $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$ can be assumed to be normalised. Linear CKA is invariant to orthogonal transformations (rotations or reflections), isotropic scaling, and translations of the representations, so it captures only the relational structure shared between the two spaces.
Representational Similarity Analysis  (Kriegeskorte et al., ).
@@
Representational Similarity Analysis (RSA) compares the geometry of representations via their Representational Dissimilarity Matrices (RDMs). For each representation, we compute pairwise dissimilarities between stimuli using 1Pearson correlation1-$\text${Pearson correlation}, yielding an $M\times M$ RDM that encodes the relational structure. Model similarity is then quantified as the Pearson correlation between their RDMs. RSA is invariant to orthogonal transformations and reflects how models structure their representational spaces, independent of the specific features they encode.
Soft Matching (Khosla & Williams, 2024).
@@
Soft Matching (SoftMatch) generalizes permutation distance (Ding et al., 2021) to representations with different numbers of units by relaxing permutations to “soft permutations.” Specifically, consider a non-negative matrix $\mathbf{P}\in\mathbb{R}^{N_{i}\times N_{j}}$ whose rows each sum to $1/N_{i}$ and columns to $1/N_{j}$. The set of such matrices defines a transportation polytope (De Loera & Kim, 2013), $\mathcal{T}(N_{i},N_{j})$. The optimization problem is
@@
Procrustes analysis finds the orthogonal transformation that best aligns two representations while preserving geometry. For unequal dimensions, the smaller representation is zero-padded. The optimization problem is $\min_{\mathbf{R}\in\mathcal{O}(N)}\|\mathbf{X}_{j}-\mathbf{X}_{i}\mathbf{R}\|_{2}^{2},$ where $\mathcal{O}(N)=\{\mathbf{R}\in\mathbb{R}^{N\times N}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}\}$. The optimal transformation $\mathbf{R}^{\star}$ is obtained via singular value decomposition. The similarity score is the mean unit-wise correlation between $\mathbf{X}_{j}$ and $\mathbf{X}_{i}\mathbf{R}^{\star}$.
Linear Predictivity (Yamins et al., 2014).
@@
Linear predictivity seeks an unconstrained linear transformation that best predicts one representation from another: $\min_{\mathbf{L}}\|\mathbf{X}_{j}-\mathbf{X}_{i}\mathbf{L}\|_{2}^{2}$. The optimal mapping $\mathbf{L}^{\star}$ is estimated via ordinary least squares. The final similarity score is the mean unit-wise correlation between $\mathbf{X}_{j}$ and $\mathbf{X}_{i}\mathbf{L}^{\star}$.
Average Baseline.
@@
To provide a baseline that naively uses all metrics’ information, we symmetrized and min-max rescaled all metrics’ result matrices and simply averaged them.
2.3 Separation Ability Metrics
We next describe the measures used to evaluate how well representational metrics capture separability between model families. Because family separation is inherently bidirectional, we compute directional scores in both directions and report the average as the final result.
Contrastive Ratio.

### v4

# Comparing and Integrating Different Notions of Representational Correspondence in Neural Systems

@@ Abstract
The extent to which different biological and artificial neural systems rely on equivalent internal representations to support similar tasks remains a central question in neuroscience and machine learning. Prior work typically compares systems using a single representational similarity metric, even though different metrics emphasize distinct facets of representational correspondence. Here we evaluate a suite of representational similarity measures by asking how well each metric recovers known structure across two domains: for artificial models, whether procedurally dissimilar models (differing in architecture or training paradigm) are assigned lower similarity than procedurally matched models; and for neural data, whether responses from distinct cortical regions are separated while responses from the same region align across subjects. Across both vision models and neural recordings, metrics that preserve representational geometry or tuning structure more reliably separate this structure than more flexible mappings such as linear predictivity. To integrate these complementary facets, we adapt Similarity Network Fusion, originally developed for multi-omics integration, to combine similarity graphs across metrics. The resulting fused similarity yields sharper separation of procedurally defined model families and, when applied to neural data, recovers a clearer hierarchical organization of the ventral visual stream that aligns more closely with established anatomical and functional hierarchies than single metrics. Overall, this approach reveals which dimensions of representational correspondence recover meaningful structure in models and brains, and how complementary notions of similarity can be integrated.
Representation Analysis, Model Typology, Vision Models, Transformer, CNN
@@ 1 Introduction
Understanding when two neural systems, biological or artificial, rely on equivalent internal representations is a central challenge in neuroscience and machine learning. Researchers typically compare representations using a single similarity metric, such as Representational Similarity Analysis (RSA) (Kriegeskorte et al., ), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), or linear predictivity (Yamins et al., 2014). Crucially, these metrics are not interchangeable. Each metric captures a distinct notion of correspondence, preserving representational geometry, aligning unit-level tuning, or measuring alignment in linearly accessible information, and therefore makes distinct assumptions about which representational transformations should be considered irrelevant. As a result, different metrics can yield qualitatively different conclusions about which systems are similar or distinct(Soni et al., 2024; Cloos et al., 2024).
This observation points to a deeper issue: representational similarity is inherently multidimensional. Some aspects of representation may be widely shared across systems, reflecting convergent computational solutions, while others may encode signatures that are specific to particular architectures, training objectives, or biological substrates. As a result, representational comparison remains methodologically fragmented, with different similarity measures often yielding conflicting views of correspondence. This raises a fundamental question: which dimensions of representational correspondence can recover meaningful structure?
Answering this question requires domain-appropriate criteria. In artificial neural networks, a natural benchmark is procedural correspondence: whether a similarity measure distinguishes models trained under different architectures or learning paradigms while grouping procedurally matched models. In neural data, an analogous benchmark is anatomical–functional organization: whether responses from distinct cortical regions are separable while responses from the same region align across subjects. Together, these criteria provide principled tests for evaluating which representational similarity measures capture meaningful structure.
Using vision models and visual cortical recordings as a controlled testbed, this work makes three central contributions. First, we provide a systematic evaluation of representational similarity measures capturing distinct dimensions of correspondence (e.g. geometry, unit-level tuning, and linearly accessible information) using principled criteria in both domains. We find that metrics preserving representational geometry or unit-level tuning (e.g., RSA and soft-matching variants (Khosla & Williams, 2024)) more reliably recover known structure, separating procedurally distinct models and differentiating cortical regions, yet more flexible mappings such as linear predictivity yield weaker discrimination.
Second, recognizing that these representational dimensions are complementary and that no single metric provides a complete characterization, we introduce an integrative framework based on Similarity Network Fusion (SNF) (Wang et al., 2014). By combining similarity graphs derived from multiple metrics, the fused representation sharpens recovery of procedural structure among models and reveals clearer anatomical–functional organization across cortical regions than any individual metric alone.
Finally, we show that integrating complementary similarity dimensions enables the construction of meaningful, data-driven typologies of both artificial models and neural systems. This approach follows empirical traditions in psychology, neuroscience, and genetics (Wang et al., 2014; Letwin et al., 2006; Echtermeyer et al., 2011; Mukamel & Ngai, 2019) where researchers identify natural groupings by combining multiple behavioral or molecular indices. For AI models, clustering the SNF-fused similarity matrix reveals groupings that emerge from how systems organize information rather than from prior assumptions about architecture or training tasks. For neural data, the fused similarities expose clear cross-subject structure: responses from the same cortical area form compact clusters that are invariant across subjects, with robust separation between areas. These results demonstrate that integrating across representational dimensions not only improves discriminability, but also supports principled discovery of organization consistent with known anatomical and functional hierarchies.
@@
Figure 1: Top left: Each representational metric defines a pairwise similarity matrix over models/brain regions. Bottom left: Each matrix is visualized as an affinity graph, with nodes representing models/brain regions and edge widths reflecting pairwise similarity strength; weak similarities below a threshold are omitted for clarity. Right: A consensus matrix obtained via Similarity Network Fusion (SNF) highlights relations consistently supported across metrics while leveraging complementary signals. In the fused graph, solid edges denote agreement across all metrics, dotted edges indicate partial support; strong but uncorroborated edges may persist with reduced weight (e.g., edge 4–5); weak AND metric-specific connections are typically suppressed (e.g., edge 1–5).
@@ 2.1 Model Selection and Dataset
To evaluate how different representational similarity measures recover known procedural structure, we curated a controlled population of vision models in which architecture and training objective vary systematically while other factors are held fixed. We define a model’s procedure as the combination of its architecture and training paradigm, and treat models sharing the same procedure as belonging to the same model family. We analyze 35 vision models across four primary categories: supervised Convolutional Neural Networks (CNNs), self-supervised CNNs, supervised Transformers, and self-supervised Transformers. We treat ConvNeXt (Liu et al., 2022) and Swin (Liu et al., 2021) as distinct families due to their hybrid nature—ConvNeXt incorporates Transformer-inspired design principles within a convolutional architecture, while Swin introduces CNN-like inductive biases into the Transformer framework. We deliberately restricted the model set to encoders pre-trained on ImageNet-1k. This design isolates the effects of architecture and training objective within a shared data and label space so that differences in training data distribution do not impact. For datasets, we use the ImageNet-1k (Deng et al., 2009) and Ecoset (Mehrer et al., 2021) validation sets and CIFAR10 (Krizhevsky, 2009) and CIFAR100 (Krizhevsky, 2009) test sets. More details are provided in Appendix A and B. For brain data, we consider fMRI responses to the 1,000 shared natural images from the Natural Scenes Dataset (NSD) (Allen et al., ), spanning 10 visual regions across 4 subjects. This dataset enables analogous population-level comparisons in the brain, allowing us to evaluate whether representational similarity measures separate responses from distinct cortical regions while preserving cross-subject alignment within the same region.
@@ 2.2 Representational Metrics
We evaluate widely used similarity metrics that differ in the flexibility of the mappings they permit—from permutation-based alignments (soft-matching) to rigid geometric transformations (Procrustes) to looser linear mappings (linear predictivity) as well as non-fitting approaches that compare representational geometry directly (RSA). Consider two representations $\mathbf{X}_{i}\in\mathbb{R}^{M\times N_{i}}$ and $\mathbf{X}_{j}\in\mathbb{R}^{M\times N_{j}}$ from different models, where $M$ denotes the number of stimuli and $N_{i},N_{j}$ denote the number of units. All representations are mean-centered along the sample dimension. For metrics requiring a fitting procedure (e.g., Soft matching, linear predictivity, Procrustes), similarity values reflect the mean 5-fold cross-validation score.
@@
Singular Vector Canonical Correlation Analysis (Raghu et al., 2017). SVCCA first applies singular value decomposition (SVD) to the representation matrices from two models or layers to isolate their most informative directions: $\mathbf{X}_{i}=\mathbf{U}_{i}\Sigma_{i}\mathbf{V}_{i}^{\top}$ and $\mathbf{X}_{j}=\mathbf{U}_{j}\Sigma_{j}\mathbf{V}_{j}^{\top}$. Retaining the top $N^{\prime}_{i}$ and $N^{\prime}_{j}$ singular vectors that explain $99\%$ of the variance yields the reduced representations: $\mathbf{X}^{\prime}_{i}=\mathbf{U}_{i}^{(N^{\prime}_{i})\!\top}\mathbf{X}_{i}$ and $\mathbf{X}^{\prime}_{j}=\mathbf{U}_{j}^{(N^{\prime}_{j})\!\top}\mathbf{X}_{j}$ whose dominant singular directions capture a disproportionate share of the total information. Canonical correlation analysis (CCA) (Hardoon et al., 2004) is then applied to these reduced matrices to find linear projections $\mathbf{A}$ and $\mathbf{B}$ that maximize their correlation: $\mathbf{Q}=\max_{\mathbf{A},\,\mathbf{B}}\operatorname{corr}\!\bigl(\mathbf{A}\mathbf{X}^{\prime}_{i},\;\mathbf{B}\mathbf{X}^{\prime}_{j}\bigr)$ subject to unit-variance constraints, providing a final similarity score that reflects how closely the subspaces of the two representations align.
@@
Projection-Weighted Canonical Correlation Analysis (Morcos et al., 2018). Compared to SVCCA, PWCCA did not apply SVD before CCA and it re-weights the canonical directions according to their contribution to the original representation. After CCA, we obtain canonical vectors $\mathbf{A}$ and $\mathbf{B}$ and their corresponding correlations $q_{i}$ for $i=1,2,\ldots,k$, where $k=\min\{N_{i},N_{j}\}$. Instead of giving each direction equal weight, PWCCA projects the unreduced representation $\mathbf{X}_{i}$ onto its own canonical vectors to measure how strongly each one reconstructs the data. The projection weight for the $i^{th}$ canonical direction is $\alpha_{i}=\frac{\lVert\mathbf{X}_{i}\mathbf{a}_{i}\rVert_{1}}{\sum_{j=1}^{k}\lVert\mathbf{X}_{i}\mathbf{a}_{j}\rVert_{1}},$ where $\mathbf{a}_{i}$ is the $i^{th}$ column of $\mathbf{A}$. The final PWCCA similarity is then the weighted sum of canonical correlations: $\sum_{i=1}^{k}\alpha_{i}q_{i}$ which emphasizes directions that explain the greatest fraction of variance in $X_{i}$ and down-weights noisy, low-variance components, yielding a more faithful measure of representational similarity.
@@
Linear Centered Kernel Alignment (Kornblith et al., 2019). Linear CKA provides a scalar measure of how similarly two sets of representations capture the relationships among the same collection of samples. It is is defined as $\frac{\|\mathbf{X}_{i}^{T}\mathbf{X}_{j}\|_{F}^{2}}{\|\mathbf{X}_{i}^{T}\mathbf{X}_{i}\|_{F}\|\mathbf{X}_{j}^{T}\mathbf{X}_{j}\|_{F}}$, where $\|.\|_{F}$ is the frobenius norm, and $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$ can be assumed to be normalised. Linear CKA is invariant to orthogonal transformations (rotations or reflections), isotropic scaling, and translations of the representations, so it captures only the relational structure shared between the two spaces.
@@
Representational Similarity Analysis  (Kriegeskorte et al., ). Representational Similarity Analysis (RSA) compares the geometry of representations via their Representational Dissimilarity Matrices (RDMs). For each representation, we compute pairwise dissimilarities between stimuli using 1Pearson correlation1-$\text${Pearson correlation}, yielding an $M\times M$ RDM that encodes the relational structure. Model similarity is then quantified as the Pearson correlation between their RDMs. RSA is invariant to orthogonal transformations and reflects how models structure their representational spaces.
@@
Soft Matching (Khosla & Williams, 2024). Soft Matching (SoftMatch) generalizes permutation distance (Ding et al., 2021) to representations with different numbers of units by relaxing permutations to “soft permutations.” Specifically, consider a non-negative matrix $\mathbf{P}\in\mathbb{R}^{N_{i}\times N_{j}}$ whose rows each sum to $1/N_{i}$ and columns to $1/N_{j}$. The set of such matrices defines a transportation polytope (De Loera & Kim, 2013), $\mathcal{T}(N_{i},N_{j})$. The optimization problem is
@@
Procrustes Alignment (Ding et al., 2021). Procrustes analysis finds the orthogonal transformation that best aligns two representations while preserving geometry. For unequal dimensions, the smaller representation is zero-padded. The optimization problem is $\min_{\mathbf{R}\in\mathcal{O}(N)}\|\mathbf{X}_{j}-\mathbf{X}_{i}\mathbf{R}\|_{2}^{2},$ where $\mathcal{O}(N)=\{\mathbf{R}\in\mathbb{R}^{N\times N}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}\}$. The optimal transformation $\mathbf{R}^{\star}$ is obtained via singular value decomposition. The similarity score is the mean unit-wise correlation between $\mathbf{X}_{j}$ and $\mathbf{X}_{i}\mathbf{R}^{\star}$.
@@
Linear Predictivity (Yamins et al., 2014). Linear predictivity seeks an unconstrained linear transformation that best predicts one representation from another: $\min_{\mathbf{L}}\|\mathbf{X}_{j}-\mathbf{X}_{i}\mathbf{L}\|_{2}^{2}$. The optimal mapping $\mathbf{L}^{\star}$ is estimated via ordinary least squares. The final similarity score is the mean unit-wise correlation between $\mathbf{X}_{j}$ and $\mathbf{X}_{i}\mathbf{L}^{\star}$.
@@
Average Baseline. To provide a baseline that naively uses all metrics’ information, we symmetrized and min-max rescaled all metrics’ matrices and simply averaged them.
2.3 Separability Metrics
We next describe the measures used to assess how well representational similarity metrics recover procedural distinctions among models (i.e. how well they separate different model families).